# BICEP/Keck XIX: Extremely Thin Composite Polymer Vacuum Windows for BICEP and Other High Throughput Millimeter Wave Telescopes

### v1

@@ 3.1 Anti-reflection
The AR coat deployed on the previous bulk polyethylene windows were generated before AR process improvements described in Eiben et al. (2022) and Dierickx et al. (2023), and therefore the designed null is off-center in the band, as seen in Figure 8. Additionally, the thin window allows for a second designed reflection null in-band, and eliminates the reflection fringes that the thick slab produces. The very thin nature of the laminate windows allows for exceptionally low reflection with a single layer AR coat. However, we must laminate at high pressure to generate a laminate window with the expected index, which is vital for designing AR coats. In the next section we discuss an additional potential optical issue with poor lamination, which may affect the polarized transfer function of the receiver.
3.2 Birefringence
The BICEP/Keck Collaboration is aiming to measure the polarization of a cosmological signal to very high precision. The B-mode angular power spectra are expected to be multiple orders of magnitude below the temperature anisotropies. Due to the highly aligned structure of the HMPE fibers within the woven fabric at macroscopic scales, there is a legitimate concern that the woven structure we are incorporating into these windows may induce a polarized response, reducing our sensitivity to primordial polarization states. An instrumental polarized effect, such as birefringence, from the window would induce correlations that our normal analysis self-calibrates for and eliminates (Keating et al., 2013; BICEP/Keck Collaboration et al., 2021). Nevertheless, we constrain any potential polarizing effect from the laminate window using both the benchtop measurements described in this section and direct measurements on the BICEP3 receiver in Section 4.1.
@@ 3.2 Birefringence
If there is a birefringent effect from the difference in index between the HMPE fibers and the surrounding LDPE, the windows are thin enough that there should not be a long enough path to produce a significant phase offset. However, it is also possible that the HMPE weave induces other polarization effects, such as polarized scatter, that in practice act like a birefringent material. Additionally, as discussed previously in Section 3.1, the lamination quality (fill factor) has an effect on the index of the laminate. The lower pressure laminates have a lower fill factor and therefore lower index; the HMPE weave in lower pressure laminates is even more visible due to the poor lamination (see Figure 3). If birefringence or other polarized effects were to occur, it is more likely to be measurable in a low pressure laminate where gaps between the HMPE and LDPE are more prevalent.
However, multi-layer laminates—such as the deployed windows—have three layers of HMPE weave rotated 30 degrees offset from each other, as discussed in Section 2.1. Multiple weave orientations may obscure the relation between the weave and the phase offset. To better understand the polarization properties of the laminated fibers, we test single layer HMPE laminates to reduce potentially confounding additional angles.
We can compare between a poorly laminated (1 atm) single layer laminate, a well laminated (10 atm) single layer laminate, and a thin extruded bulk HDPE sample. Extruded bulk HDPE windows have been observed to exhibit birefringent properties at far IR frequencies, so we use it as a detection upper limit (Paine & Turner, 2013). To measure the transmission () phase we used the previously described VNA set up with an additional receive-only port using an even harmonic mixer to downconvert the received signal. This setup is shown in the left of Figure 9. When we introduce a material to the beam of our VNA, it will introduce a phase offset due to the change in electrical length of the path between the horns that varies as:
where is the phase offset (in radians), is the index of refraction, is the thickness of the material, and is the wavelength of the incident light. The electrical length is the product of the index and the thickness of the material. Therefore, changes in the index at different orientations will result in a different offset of phase, and a different electrical length.
@@
We fit both a flat line and a sin(2) wave to the measurements to test the null hypothesis. The sine wave fit has two more degrees of freedom than the flat line, and therefore the between the flat fits and the sin(2) fits would have to exceed the extra degrees of freedom to be significant, which they do not.
We might expect more than one fast or slow axis induced by the HMPE weave in the laminate. However, these measurements cannot constrain more than a uniaxial birefringent effect. We would also expect that extruded HDPE would act more like a traditional uniaxial birefringent material—as we suspect that the extrusion should induce a preferred orientation of polymer crystals along one axis in the manufacturing process—which should be captured by these measurements (Paine & Turner, 2013). The HDPE was also the thickest of the samples tested, and therefore would induce the largest phase difference and the greatest chance of detection. There is no strong evidence for a detection of birefringence in the extruded HDPE sample. As such, we set an upper limit on birefringence in the laminate windows at the maximum level of the fit sin(2) amplitudes. The maximum fit amplitude in electrical length is approximately 0.008 mm for any of the materials, or approximately a 0.02 change in index for the thinnest laminate.
An upper limit change of 0.02 in the index is supported by the typical repeatability of index of refraction fits to laminate windows (fit indexes range between 1.536 to 1.557, as discussed in Section 3.1). If we calculate the impact of such a birefringent window on a fully polarized incident Q beam, we can put constraints on those potential impacts. We can take the combination of window parameters that will produce the worst phase retardation: the thickest deployed window (1.4 mm on BICEP3), and the highest frequency a deployed window is designed for (170 GHz, the upper edge of the BA150 band) to constrain the worst case scenario. The window can be modeled with a Mueller matrix such as:
@@
The maximum reduction of polarization efficiency is estimated by the difference of the full intensity to the out Q parameter:
Which we can see remains very low (0.3%), roughly half of the previously constrained BICEP3 polarization efficiency (BICEP/Keck Collaboration et al., 2022). Fully characterizing the polarization efficiency of a receiver requires a full end-to-end polarization measurement, which is very difficult to accurately conduct (Cornelison et al., 2022; BICEP/Keck Collaboration et al., 2024). We plan to run another absolute polarization calibration on BICEP3 with a thin laminate window in the upcoming (2024–2025) summer season, and the change in polarization angle of detectors will then be directly compared.
Though a hypothetically birefringent window results in negligibly uniform rotations of polarized light to first order, a higher order change in the polarized transfer function of the telescope would not be well characterized by these estimates. In Section 4.1 we explore the effects of how the thin window changed the residual beams in BICEP3, and therefore the estimated temperature to polarization leakage.
4 Optical Characterization on BICEP3
@@
There were three major sources of concern regarding the impact of this new window on the optical performance of BICEP3. The polarized residual beam response, also known as temperature-to-polarization leakage, is discussed in Section 4.1. The optical load from the window seen by the detectors is discussed in Section 4.2. Finally, the measured change in the detector white noise within the receiver is discussed in Section 4.3.
4.1 Temperature-to-Polarization Leakage
As discussed in Section 3.2, the polarized transfer function of the telescope may be changed by the thin window. Since we carry out polarimetry by differencing co-located, orthogonally-polarized detectors, any deviations in beam response between orthogonal pairs results in differential beam response. By doing a Taylor expansion of the differential beam, we can predict the shape of differential beam response for the lowest order modes (up to order 2). We then regress our maps against a known temperature sky (i.e., a Planck temperature map) to fit the amplitude of these modes, and deproject their response (Planck Collaboration et al., 2020b). The residual beam response after deprojection is responsible for temperature-to-polarization (T-to-P) leakage, as the undeprojected beam response leaks the bright temperature sky into our polarization maps. This is currently our main source of systematic uncertainty (BICEP2 Collaboration et al., 2019; BICEP/Keck Collaboration et al., 2021).
@@ 4.1 Temperature-to-Polarization Leakage
4.2 Measured Optical Load
The anticipated decrease in instrumental optical load from decreasing the window thickness is a primary driver for developing thin laminate windows, as shown in Figure 1. We can test the difference in optical load with our receivers themselves, as they are highly sensitive and offer the most direct measure of receiver response. We call such measurements “radiometric,” as they directly measure the change in optical load with the receiver.
@@ 4.2 Measured Optical Load
As expected, the thin 1.4 mm HMPE window causes significantly less optical loading on the detectors compared to the 31.8 mm (1.25”) UHMWPE window. However, the magnitude of the difference is not as large as expected from the assumption that the polyethylene absorption losses are the same: if the loss tangent is the same in both polyethylene windows the difference in thickness between windows would result in a factor of 20 difference in optical power. With the radiometric measurements we only measure a factor of 6.6 difference in optical loading. We therefore conclude that the two polyethylenes likely have different loss tangents. This is supported by the differences in noise equivalent temperature (NET) between years, as discussed in Section 4.3.
@@
As previously described in Section 2, the forebaffle (shown in Figure 2) terminates scattered light exiting the receiver on an absorber of known temperature. By taking PLC measurements with the materials with and without the forebaffles on the receiver, we can directly measure the excess in optical load associated with the scattered light through these windows. Shown in the right side of Figure 10 the measured scattered power associated with the window materials is very small, and for both materials within two standard deviations of zero. These measurements put upper limits on the scattered fraction from these materials, at roughly 4% for the UHMWPE and 10% for the HMPE laminate. The amplitude of these upper limits on scattered power from window material have been corroborated by benchtop scatterometer measurements reported in Corrigan (2019).
@@
4.3 Noise Equivalent Temperature
@@
Between 2018 and 2019: A window change. The window thickness decreased (31.8 mm to 24.6 mm), and the material changed (HDPE to UHMWPE). The median per detector yearly NET increased 1.7% (263 to 268 ).
@@
Between 2021 and 2022: High optics temperatures. A leaky flange led to many cryogenic optics’ temperatures to be elevated. The median yearly NET increased 2.7% (269 to 276 ).
@@
Between 2022 and 2023: A window change. The window thickness decreased (24.6 mm to 1.4 mm), and the material changed (UHMWPE to HMPE laminate). The median per detector yearly NET decreased 8.6% (276 to 252 ).
@@
A reduction in the white noise of a receiver directly corresponds to a reduction of integration time to reach a target sensitivity. The integration time of the receiver scales as ; an improvement of noise in BICEP3 of 6% (from 268 K in 2019 thru 2021 to 252 K in 2023) allows the receiver to map approximately 13% faster. At higher frequencies, because of the greater fractional load reduction from thin windows, we expect this improvement in mapping speed to be even larger. At a band center of 150 GHz, a thin 1.4 mm window should decrease the NET 12% compared to a 24.6 mm slab with the same loss tangent (from 319 K to 281 K), resulting in a 22% increase in mapping speed. We are currently analyzing the noise data from BA150 (deployed with a thin window in 2022) to confirm this expected improvement. The highest frequency bands will have an even greater improvement: at band centers 220 and 270 GHz we expect a 14% change in NET (575 K from 668 K and 1073 K from 1251 K, respectively), and 26% faster mapping speed. When BA220/270 deploys and begins observing in the coming austral summer, we also anticipate this reduction in noise.
5 Conclusion
Laminated high modulus polyethylene windows provide a unique opportunity to reduce instrumental loading on millimeter wavelength telescopes. The reflection loss of the thin window can be reduced over a broad bandwidth by tuning the thickness of the window itself, providing a rare opportunity to reduce the net reflectivity of the window to less than 0.1% with single layer anti-reflection coats. Birefringence through the HMPE window is shown to be negligible.
We have performed extensively mechanical tests on HMPE laminates. The mechanical safety factors of the thin windows are at least 5.7 at sea level, and 8.2 at operating altitude. Though the windows do deflect significantly, they are not expected to creep into contact with the filters below them within the next five years of continuous use.
@@ 5 Conclusion
We plan to deploy thin windows on the next two CMB receivers to be sent to the South Pole; on the high frequency BA receiver (220/270 GHz) in 2024 and PreSAT (95/155 GHz) in a subsequent season (The BICEP/Keck Collaboration et al., 2024; Petroff et al., 2024). The primary challenge remaining for those HMPE windows are their anti-reflection coats, partially due to the large fractional bandwidths of each receiver (0.4 for BA 220/270, 0.7 for PreSAT) (Eiben et al., 2024). A similarly difficult AR coated HMPE window will be designed for prototype wSMA receivers, covering 190 to 380 GHz, though it will require a much smaller diameter window than modern CMB small aperture telescopes (Grimes et al., 2024).
@@ 6 Acknowledgments
We thank the staff of the U.S. Antarctic Program and the South Pole Station who have enabled this research. We thank the BICEP3 winter-overs Sam Harrison (2015), Hans Boenish (2016, 2018), Grantland Hall (2017), Ta-Lee Shue (2019), Paula Crock (2020), Calvin Tsai (2021), Karsten Look (2022), and Manwei Chan (2023) as well as the BICEP Array winter-overs Thomas Leps (2022) and Anthony DeCicco (2023) for their invaluable work maintaining our telescopes throughout the winter observing season. We also thank Dr. Marion Dierickx for her significant contributions during her tenure as the BICEP Operations Manager.
We also thank Vedant Chandra, Victoria DiTomasso, Christina Eiben, Dagmar Eiben, Amanda MacLean, Kevin Ortiz Ceballos, Locke Patton, Andrew Saydjari, and Turner Woody for their generous feedback over the years of work on this paper.

### v2

@@ III.1 Anti-reflection
The AR coat deployed on the previous bulk polyethylene windows were generated before AR process improvements described in Eiben et al. (2022) and Dierickx et al. (2023), and therefore the designed null is off-center in the band, as seen in Figure 8. Additionally, the thin window allows for a second designed reflection null in-band, and eliminates the reflection fringes that the thick slab produces. The very thin nature of the laminate windows allows for exceptionally low reflection with a single layer AR coat. However, we must laminate at high pressure (above 4 atm) to generate a laminate window with the expected index, which is vital for designing AR coats. In the next section we discuss an additional potential optical issue with poor lamination, which may affect the polarized transfer function of the receiver.
III.2 Birefringence
The BICEP/Keck Collaboration is aiming to measure the polarization of a cosmological signal to very high precision. The B-mode angular power spectra are expected to be multiple orders of magnitude below the temperature anisotropies. Due to the highly aligned structure of the HMPE fibers within the woven fabric at macroscopic scales, there is a legitimate concern that the woven structure we are incorporating into these windows may induce a polarized response, reducing our sensitivity to primordial polarization states. An instrumental polarized effect, such as birefringence, from the window would induce correlations that our normal analysis self-calibrates for and eliminates (Keating et al., 2013; BICEP/Keck Collaboration et al., 2021). Nevertheless, we constrain any potential polarizing effect from the laminate window using both the benchtop measurements described in this section and direct measurements on the BICEP3 receiver in Section IV.1.
@@ III.2 Birefringence
If there is a birefringent effect from the difference in index between the HMPE fibers and the surrounding LDPE, the windows are thin enough that there should not be a long enough path to produce a significant phase offset. However, it is also possible that the HMPE weave induces other polarization effects, such as polarized scatter, that in practice act like a birefringent material. Additionally, as discussed previously in Section III.1, the lamination quality (fill factor) has an effect on the index of the laminate. The lower pressure laminates have a lower fill factor and therefore lower index; the HMPE weave in lower pressure laminates is even more visible due to the poor lamination (see Figure 3). If birefringence or other polarized effects were to occur, it is more likely to be measurable in a low pressure laminate where gaps between the HMPE and LDPE are more prevalent.
However, multi-layer laminates—such as the deployed windows—have three layers of HMPE weave rotated 30 degrees offset from each other, as discussed in Section II.1. Multiple weave orientations may obscure the relation between the weave and the phase offset. To better understand the polarization properties of the laminated fibers, we test single layer HMPE laminates to reduce potentially confounding additional angles.
We can compare between a poorly laminated (1 atm) single layer laminate, a well laminated (10 atm) single layer laminate, and a thin extruded bulk HDPE sample. Extruded bulk HDPE windows have been observed to exhibit birefringent properties at far IR frequencies, so we use it as a detection upper limit (Paine and Turner, 2013). To measure the transmission () phase we used the previously described VNA set up with an additional receive-only port using an even harmonic mixer to downconvert the received signal. This setup is shown in the left of Figure 9. When we introduce a material to the beam of our VNA, it will introduce a phase offset due to the change in electrical length of the path between the horns that varies as:
where is the change in phase (in radians) referenced against the same length without the material (), is the index of refraction, is the thickness of the material, and is the wavelength of the incident light. The electrical length is the product of the index and the thickness of the material. Therefore, changes in the index at different orientations will result in a different offset of phase, and a different electrical length.
@@
We fit both a flat line and a sin(2) wave to the measurements to test the null hypothesis. The sine wave fit has two more degrees of freedom than the flat line, and therefore the between the flat fits and the sin(2) fits would have to exceed the extra degrees of freedom to be significant. The between each materials’ fits are approximately 0.6 (specifically, the HDPE , the 1 atm HMPE and the 10 atm HMPE )—well below the two extra degrees of freedom in the birefringent sin(2) model—which does not allow us to reject the null hypothesis.
We might expect more than one fast or slow axis induced by the HMPE weave in the laminate. However, these measurements cannot constrain more than a uniaxial birefringent effect. We would also expect that extruded HDPE would act more like a traditional uniaxial birefringent material—as we suspect that the extrusion should induce a preferred orientation of polymer crystals along one axis in the manufacturing process—which should be captured by these measurements (Paine and Turner, 2013). The HDPE was also the thickest of the samples tested, and therefore would induce the largest phase difference and the greatest chance of detection. There is no strong evidence for a detection of birefringence in the extruded HDPE sample. As such, we set an upper limit on birefringence in the laminate windows at the maximum level of the fit sin(2) amplitudes. The maximum fit amplitude in electrical length is approximately 0.008 mm for any of the materials, or approximately a 0.02 change in index for the thinnest laminate.
An upper limit change of 0.02 in the index is supported by the typical repeatability of index of refraction fits to laminate windows (fit indexes range between 1.536 to 1.557, as discussed in Section III.1). If we calculate the impact of such a birefringent window on a fully polarized incident Q beam, we can put constraints on those potential impacts. We can take the combination of window parameters that will produce the worst phase retardation: the thickest deployed window (1.4 mm on BICEP3), and the highest frequency a deployed window is designed for (170 GHz, the upper edge of the BA150 band) to constrain the worst case scenario. The window can be modeled with a Mueller matrix such as:
@@
The maximum reduction of polarization efficiency is estimated by the difference of the full intensity to and the output Q parameter:
Which we can see remains very low (0.3%), roughly half of the previously constrained BICEP3 polarization efficiency (BICEP/Keck Collaboration et al., 2022). Fully characterizing the polarization efficiency of a receiver requires a full end-to-end polarization measurement, which is very difficult to accurately conduct (Cornelison et al., 2022; BICEP/Keck Collaboration et al., 2024). We plan to run another absolute polarization calibration on BICEP3 with a thin laminate window in an upcoming summer season, and the relative change in polarization angle of detectors will then be directly compared.
Though a hypothetically birefringent window results in negligibly uniform rotations of polarized light to first order, a higher order change in the polarized transfer function of the telescope would not be well characterized by these estimates. In Section IV.1 we explore the effects of how the thin window changed the residual beams in BICEP3, and therefore the estimated temperature to polarization leakage.
IV Optical Characterization on BICEP3
@@
There were three major sources of concern regarding the impact of this new window on the optical performance of BICEP3. The polarized residual beam response, also known as temperature-to-polarization leakage, is discussed in Section IV.1. The optical load from the window seen by the detectors is discussed in Section IV.2. Finally, the measured change in the detector white noise within the receiver is discussed in Section IV.3.
IV.1 Temperature-to-Polarization Leakage
As discussed in Section III.2, the polarized transfer function of the telescope may be changed by the thin window. Since we carry out polarimetry by differencing co-located, orthogonally-polarized detectors, any deviations in beam response between orthogonal pairs results in differential beam response. By doing a Taylor expansion of the differential beam, we can predict the shape of the differential beam response for the lowest order modes (up to order 2). We then regress our maps against a known temperature sky (i.e., a Planck temperature map) to fit the amplitude of these modes, and deproject their response (Planck Collaboration et al., 2020b). The residual beam response after deprojection is responsible for temperature-to-polarization (T-to-P) leakage, as the undeprojected beam response leaks the bright temperature sky into our polarization maps. This is currently our main source of systematic uncertainty (BICEP2 Collaboration et al., 2019; BICEP/Keck Collaboration et al., 2021).
@@ IV.1 Temperature-to-Polarization Leakage
IV.2 Measured Optical Load
The anticipated decrease in in-band instrumental optical load from decreasing the window thickness is a primary driver for developing thin laminate windows, as shown in Figure 1. We can test the difference in optical load with our receivers themselves, as they are highly sensitive and offer the most direct measure of receiver response. We call such measurements “radiometric,” as they directly measure the change in optical load with the receiver.
@@ IV.2 Measured Optical Load
As expected, the thin 1.4 mm HMPE window causes significantly less optical loading on the detectors compared to the 31.8 mm (1.25”) UHMWPE window. However, the magnitude of the difference is not as large as expected from the assumption that the polyethylene absorption losses are the same: if the loss tangent is the same in both polyethylene windows the difference in thickness between windows would result in a factor of 20 difference in optical power. With the radiometric measurements we only measure a factor of 6.6 difference in optical loading. We therefore conclude that the two polyethylenes likely have different loss tangents. This is supported by the differences in noise equivalent temperature (NET) between years, as discussed in Section IV.3.
@@
As previously described in Section II, the forebaffle (shown in Figure 2) terminates scattered light exiting the receiver on an absorber of known temperature. By taking PLC measurements with the materials with and without the forebaffles on the receiver, we can directly measure the excess in optical load associated with the scattered light through these windows. Shown in the right side of Figure 10 the measured scattered power associated with the window materials is very small, and for both materials within two standard deviations of zero. These measurements put upper limits on the scattered fraction from these materials, at roughly 4% for the UHMWPE and 10% for the HMPE laminate. The amplitude of these upper limits on scattered power from window material have been corroborated by benchtop scatterometer measurements reported in Corrigan (2019).
@@
IV.3 Noise Equivalent Temperature
@@
Between 2018 and 2019: A window change. The window thickness decreased (31.8 mm to 24.6 mm), and the material changed (HDPE to UHMWPE). The median per detector yearly NET increased 1.9% (262 to 267 ).
@@
Between 2021 and 2022: High optics temperatures. A leaky flange led to many cryogenic optics’ temperatures to be elevated. The median yearly NET increased 3.0% (267 to 275 ).
@@
Between 2022 and 2023: A window change. The window thickness decreased (24.6 mm to 1.4 mm), and the material changed (UHMWPE to HMPE laminate). The median per detector yearly NET decreased 9.1% (275 to 252 ).
@@
A reduction in the white noise of a receiver directly corresponds to a reduction of integration time to reach a target sensitivity. The integration time of the receiver scales as ; an improvement of noise in BICEP3 of 6% (from 267 K in 2019 thru 2021 to 252 K in 2023 and 2024) allows the receiver to map approximately 13% faster. At higher frequencies, because of the greater fractional load reduction from thin windows, we expect this improvement in mapping speed to be even larger. At a band center of 150 GHz, a thin 1.4 mm window should decrease the NET 12% compared to a 24.6 mm slab with the same loss tangent (from 319 K to 281 K), resulting in a 22% increase in mapping speed. We are currently analyzing the noise data from BA150 (deployed with a thin window in 2022) to confirm this expected improvement. The highest frequency bands will have an even greater improvement: at band centers 220 and 270 GHz we expect a 14% change in NET (575 K from 668 K and 1073 K from 1251 K, respectively), and 26% faster mapping speed. When BA220/270 deploys and begins observing in the coming austral summer, we also anticipate this reduction in noise.
V Conclusion
Laminated high modulus polyethylene windows provide a unique opportunity to reduce instrumental loading on millimeter wavelength telescopes. The reflection loss of the thin window can be reduced over a broad bandwidth by tuning the thickness of the window itself, providing a valuable opportunity to reduce the net reflectivity of the window to less than 0.1% with single layer anti-reflection coats. Birefringence through the HMPE window is shown to be negligible.
We have performed extensive mechanical tests on HMPE laminates. The mechanical safety factors of the thin windows are at least 5.7 at sea level, and 8.2 at operating altitude. Though the windows do deflect significantly, they are not expected to creep into contact with the filters below them within the next five years of continuous use.
@@ V Conclusion
We will deploy thin windows on the next two CMB receivers sent to the South Pole; the window on the high frequency BA receiver (220/270 GHz) was deployed in 2024 and the mid-frequency BA receiver (95/155 GHz) in a subsequent season (The BICEP/Keck Collaboration et al., 2024; Petroff et al., 2024). The primary challenge remaining for these HMPE windows are their anti-reflection coats, partially due to the large fractional bandwidths of each receiver (0.4 for BA 220/270, 0.7 for BA 95/150) (Eiben et al., 2024). A similarly difficult AR coated HMPE window will be designed for prototype wSMA receivers, covering 190 to 380 GHz, though it will require a much smaller diameter window than modern CMB small aperture telescopes (Grimes et al., 2024).
@@ VI Acknowledgments
We thank the staff of the U.S. Antarctic Program and the South Pole Station who have enabled this research. We thank the BICEP3 winter-overs Sam Harrison (2015), Hans Boenish (2016, 2018), Grantland Hall (2017), Ta-Lee Shue (2019), Paula Crock (2020), Calvin Tsai (2021), Karsten Look (2022), Manwei Chan (2023), and Danielle Simmons (2024) as well as the BICEP Array winter-overs Thomas Leps (2022), Anthony DeCicco (2023) and Thibault Romand (2024) for their invaluable work maintaining our telescopes throughout the winter observing season. We thank Marion Dierickx for her significant contributions during her tenure as the BICEP Operations Manager.
We also thank Vedant Chandra, Victoria DiTomasso, Christina Eiben, Dagmar Eiben, Amanda MacLean, Kevin Ortiz Ceballos, Locke Patton, Andrew Saydjari, and Rebecca Woody for their generous feedback over the years of work described in this paper.